OpenStudy (anonymous):

Solve exactly in terms of ln or log: 2^2x + 2^x - 12 = 0

myininaya (myininaya):

let u=2^x then u^2=2^(2x) so you have u^2+u-12=0 (u+4)(u-3)=0 u=-4, u=3 but remember u=2^x 2^x=-4, 2^x=3 but you throw out 2^x=-4 because when you take natural log of both sides you get natural log of a negative which does not exist so just solve 2^x=3 and you are done